An airplane is flying horizontally at an altitude of 325m and velocity of 270m/s what distance before the drop target should we drop the package?
1
Expert's answer
2020-04-26T18:50:47-0400

The equations of motion of the package are "y(t) = h - \frac{g t^2}{2}", "x(t) = v_0 t", where "h" is the initial altitude, "v_0" - initial velocity of the package.

Equating altitude to zero, one can find the time "T" it will take to reach the ground: "0 = h - \frac{g T^2}{2} \Rightarrow T = \sqrt{\frac{2 h}{g}}".

Therefore, the horizontal distance before dropping the package should be "L= x(T) = v_0 T = v_0 \sqrt{\frac{2 h}{g}} \approx 2197.8 m".